this is a differential equation question


please solve them with steps and details also, make sure it is correct

1. (3 points each) Find the following functions, (a) C{(t - 1) step(t - 4)} (b) C-{ S-2 } 2s +1 (c) C-{ s2 + 6s + 25
Answer #1

Ans. a L{t-1) -1² step (t (t-w? Heue this can be written as Lfct-12 ult-4) Now Heue we know that the Heue Lfelt-a) uſt-as} ceNow het-a) = g(+) hlt-eta) getta) h(t) = g(t+a) Now put the value in equation ® e-as e L{getras Now come to the main Question-) eto Now al peu equation 20 we get (tts) u(t-l) =) els 2 (45 t 4eet+2) { teſ + L{o} + L{2s e-45 키 2, +oL{t} + + 2S {1} S3 e6 tte} S-2 Heue we know that the t{eet} - Htt-a) f(t-e) where H(t) is a Heaviside step function and heue a=4 - {e (} .td-)H(t-u) L 2 [ feruo (5-2)} Lde hay 2(t-4) H(4-4) e Ans. R I 25+1 82 + 65 +25 Now this can be written as 2 s tl 2 S + 6-5 = S2Now اتا {ation Now By the lineauity propelity of inveuse laplace transform given t{a.fun a. ft) + b. gro } - a { fro} + . Cso =) 2 니 S +3 (5+3)2 + (4) = ) 2. Terst cos(ht z) 2 e 3t cos (4t) Now solve S. (5 +372 +16 ایده ) 5 to Constantin (5+3)2 +16Now solve [Antro] t{ outle] = ) Heue L{ einutta se wa (5>0) = ) Helle w2=16 - 4 Now w S2+02 (1) s² + +42 [ حم 4 La port(tNow put all the values in equation 2 Jo We get 네 2 S+3 (S+3) + G (+3) + Ic }-{ srfant U 2. 드 S+3 (G+3) + 1G = 2e -3t os(4t) 3